# FlexQ: Efficient Post-training INT6 Quantization for LLM Serving via Algorithm-System Co-Design

### v1

@@ Abstract
In this paper, we propose FlexQ, a novel post-training INT6 quantization framework combining algorithmic innovation with system-level optimizations. FlexQ employs uniform 6-bit weight quantization across all layers, with adaptive retention of 8-bit activations in layers identified through layer-wise sensitivity analysis. To maximize hardware efficiency, we develop a specialized high-performance GPU kernel supporting matrix multiplication for W6A6 and W6A8 representations via Binary Tensor Core (BTC) equivalents, effectively bypassing the lack of native INT6 tensor cores. Evaluations on LLaMA models show FlexQ maintains near-FP16 accuracy, with perplexity increases of no more than 0.05. The proposed kernel achieves an average 1.39× speedup over ABQ-LLM on LLaMA-2-70B linear layers. End-to-end, FlexQ delivers 1.33× inference acceleration and 1.21× memory savings over SmoothQuant. Code is released at https://github.com/FlyFoxPlayer/FlexQ.
@@ 1 Introduction
  • We perform a comprehensive evaluation on the LLaMA family and demonstrate that FlexQ achieves near-FP16 perplexity (with an increase of at most 0.05) on WikiText2, indicating negligible accuracy degradation. Our kernel significantly accelerates inference, delivering a 1.39× speedup over existing approaches like ABQ-LLM. Furthermore, end-to-end evaluations reveal that FlexQ surpasses state-of-the-art methods like SmoothQuant, delivering 1.33× faster inference and 1.21× greater memory savings.
@@ 2.1 Quantization of Large Language Models
Although LLMs have demonstrated remarkable performance, their massive size poses significant deployment challenges. Model quantization is a widely adopted technique to address these issues by producing more compact model representations. The primary focus of quantization is on the weights of linear layers (i.e., matrix multiplication), which account for approximately 99% of the total model weights (Lin et al. 2024c). Activations can also be quantized during inference to further improve efficiency (Nagel et al. 2021). Quantization schemes can be categorized into symmetric and asymmetric approaches based on whether a zero point is employed to accommodate data distribution asymmetry.(Dettmers et al. 2022; Zhao et al. 2024; Liu and Yu 2025) Asymmetric quantization incorporates zero-point offsets to better capture the dynamic range of the data, thereby offering superior numerical precision. In contrast, symmetric quantization eliminates the zero-point, simplifying computations and making it more hardware-efficient. In this paper, we adopt symmetric quantization for both weights and activations. We denote the precision configuration using the notation ’WxAy’, where x/y indicates the bit-width for weights and activations, respectively.
@@ 2.2 NVIDIA GPU Architecture and Tensor Core
GPU Execution Hierarchy. NVIDIA GPUs are optimized for large-scale parallel computing, comprising multiple Streaming Multiprocessors (SMs), each integrating CUDA cores, Tensor Cores (TCs), and other specialized functional units. The execution model employs a hierarchical thread organization: individual threads serve as the fundamental execution units, with 32 threads forming a warp that executes instructions synchronously in lockstep (NVIDIA 2024). Multiple warps constitute a thread block, which shares on-chip memory resources. When launching a kernel, the configuration of the number of thread blocks and threads per block must be specified. Each thread block is assigned to a specific SM for execution, where threads within the same warp operate synchronously under the Single Instruction, Multiple Threads (SIMT) execution model.
@@ 2.3 Related Work
Weight-only Quantization. Weight-only quantization approximates weight matrices with low-bit representations to reduce computational complexity and memory footprint. GPTQ (Frantar et al. 2022) utilizes 4-bit integers combined with Hessian-based error compensation to minimize quantization errors in LLMs. AWQ (Lin et al. 2024b) and OWQ (Lee et al. 2024) emphasize the importance of quantizing weights associated with higher-magnitude activations, significantly improving quantized model performance. QuIP (Chee et al. 2023) and QuIP# (Tseng et al. 2024) achieve 2-bit quantization by employing learnable codebooks and additional fine-tuning, often coupled with vector quantization of weights. QuaRot (Ashkboos et al. 2024) implements INT4 quantization using Hadamard transformations to effectively handle outliers in linear and attention layers. Recently, Quant-LLM (Xia et al. 2024) proposes the first GPU kernel designed specifically for FP6 weight quantization. In contrast, our work applies INT6 quantization to LLM weights.
@@ 4.1 Quantization Scheme
Figure 4: Perplexity results (lower is better) for quantizing various linear layers in the LLaMA-7B and LLaMA-2-7B models on the Wikitext2 dataset.
@@
Figure 5: Overview of the FlexQ workflow on LLaMA, where ReQuant and DeQuant denote online quantization and dequantization operations, respectively, and BitPacking represents the online Bit-level Data Packing operation for activations.
@@ 4.3 FlexQ-W6Ax: Bit-Level Kernel Design
Specifically, for our GEMM kernel, given a weight matrix W and an activation matrix in bit-level packed form for computation within a single thread block, we aim to optimize the memory access pattern to enable coalescing. Each thread performs load operations with a granularity of 16 bytes (128 bits). Consequently, in a single transaction, all threads (totaling threads_num) collectively transfer 16*threads_num bytes, equivalent to threads_num INT4 elements. By applying bit-level data packing to both W and X, the data can be aligned according to structures such as p*chunk_n*chunk_k or q*chunk_m*chunk_k. This alignment ensures that memory accesses are contiguous across threads, thus facilitating efficient coalesced memory transactions.
@@
Taking the mma.m8n8k128 shape as an illustrative example, the compiler partitions each warp into eight thread groups, each comprising four consecutive threads. During data loading, the four threads within each group cooperatively load segments of the weight and activation matrices along the MMA_K dimension. Concurrently, the thread groups collaboratively handle data transfer along the [p, chunk_n] and [q, chunk_m] dimensions. Within a single iteration, each warp executes WARP_M_TILES*WARP_N_TILES BMMA operations, processing a data volume proportional to q*chunk_m*p*chunk_n*MMA_K. The intermediate results from these operations are accumulated to produce the non-weighted GEMM output.
FlexQ employs fine-grained group-wise quantization for both weights and activations. Each BMMA operation generates an intermediate result representing the multiplication of one weight group and one activation group, each associated with independently maintained quantization parameters (w_scale and x_scale). To eliminate the overhead associated with separate dequantization passes and additional memory transfers, FlexQ integrates dequantization directly into the accumulation stage of the intermediate results. Specifically, the quantization parameters w_scale and x_scale are stored as FP16 types, each thread loads the corresponding scales, which are then multiplied using the efficient half-precision parallel instruction __hmul2. The resulting dequantized intermediate values are summed across threads to produce the final FP16 result. Furthermore, by extending computations across the bit dimension among thread groups, FlexQ reduces redundant calculations, particularly in scenarios where the batch size is less than eight, thereby further optimizing GPU hardware utilization.
@@ 5.1 Experimental Setup
Dataset #Bit Method 1-7B 1-13B 1-30B 1-65B 2-7B 2-13B 2-70B
WikiText2 FP16 - 5.68 5.09 4.10 3.53 5.47 4.88 3.31
W6A6 SmoothQuant 6.03 5.42 4.55 3.88 6.20 5.18 3.69
OmniQuant 5.96 5.28 4.38 3.75 5.87 5.14 3.71
I-LLM 5.84 5.23 4.32 - 5.68 5.10 -
DuQuant 5.74 5.13 4.15 3.60 5.53 4.92 3.35
FlexQ 5.70 5.12 4.13 3.55 5.52 4.91 3.34
C4 FP16 - 7.08 6.61 5.98 5.62 6.97 6.46 5.52
W6A6 SmoothQuant 7.47 6.97 6.34 5.99 7.76 6.76 5.88
OmniQuant 7.43 6.84 6.22 5.82 7.48 6.74 5.91
I-LLM 7.32 6.79 6.25 - 7.27 6.74 -
DuQuant 7.13 6.64 6.01 5.64 7.03 6.50 5.54
FlexQ 7.11 6.63 6.00 5.63 7.02 6.49 5.53
Table 2: Perplexity results of 6-bit weight and activation quantization applied to LLaMA and LLaMA-2 models evaluated on the WikiText-2 and C4 datasets.
@@
For the accuracy experiments, we compare our method against SmoothQuant (Xiao et al. 2023), OmniQuant (Shao et al. 2024), I-LLM (Hu et al. 2024), and DuQuant (Lin et al. 2024a). Specifically, for DuQuant, we adopt the DuQuant+LWC quantization model for evaluation. To assess the efficiency of our FlexQ kernels, we compare the W6Ax variant with cuBLAS (W8A8) and ABQ-LLM (Zeng et al. 2025)’s corresponding precision kernels across various GEMM workloads on LLaMA (7B, 30B) (Touvron et al. 2023a) and LLaMA-2 (13B, 70B) models (Touvron et al. 2023b), as well as the OPT-30B model (Zhang et al. 2022). For end-to-end evaluation, we integrate our approach into FasterTransformer, comparing it against the FP16 baseline, FasterTransformer’s weight-only quantized CUTLASS (W8A16), and the SmoothQuant W8A8 implementation. It is important to note that ABQ-LLM’s quantization strategy primarily targets extremely low-bit quantization, such as INT2. Since our focus is on INT6 quantization, we exclude ABQ-LLM from the end-to-end evaluation as a baseline.
@@
We evaluate FlexQ on several LLMs: LLaMA (7B–65B), LLaMA-2 (7B–70B), and OPT (6.7B–30B). Consistent with prior work (Xiao et al. 2023; Shao et al. 2024; Hu et al. 2024; Lin et al. 2024a), we report perplexity scores on language modeling benchmarks, specifically the WikiText2 (Merity et al. 2017) and C4 (Raffel et al. 2020) datasets. To evaluate zero-shot learning performance, we adopt a benchmarking methodology similar to previous studies (Frantar et al. 2022; Xiao et al. 2023; Lin et al. 2024c; Liu et al. 2025), conducting evaluations on several standard benchmarks utilizing the lm_eval harness (Gao et al. 2021), including PIQA (Bisk et al. 2020), ARC (Clark et al. 2018), BoolQ (Clark et al. 2019), HellaSwag (Zellers et al. 2019), and Winogrande (Sakaguchi et al. 2021).
@@
Table 4: Zero-shot common-sense QA performance of LLaMA2 models using 6-bit weight and activation quantization.
@@ 5.2 Accuracy Evaluation
Table 2 reports the perplexity results of FlexQ and leading baselines on the LLaMA family. FlexQ delivers competitive results on both WikiText-2 and C4, consistently outperforming SmoothQuant, OmniQuant, and I-LLM across all model sizes. Relative to the current state-of-the-art DuQuant, FlexQ is either on par or superior, trimming perplexity by 0.01–0.05. Remarkably, FlexQ achieves nearly identical performance to FP16 baselines while being calibration-free, offering an almost lossless and highly efficient deployment solution for LLMs.
@@
To further validate our method, we compare zero-shot learning accuracy of FlexQ against selected baselines across six common-sense tasks, as shown in Tables 3 and 4. FlexQ outperforms the baselines on average across all LLaMA models of different scales. Remarkably, for certain tasks, FlexQ even surpasses the accuracy of the FP16 baseline, highlighting its ability to preserve the generalization capabilities of LLM post-quantization.
@@
Table 5: Perplexity results for LLaMA and LLaMA-2 models with 6-bit weight and activation quantization on the WikiText2 (denoted as Wiki) and C4 datasets.
@@
Figure 11: Inference latency (top) and memory utilization (bottom) measurements for the FastTransformer implementation, evaluated on an NVIDIA A6000-48GB GPU across different sequence lengths, with a comparative analysis at a fixed input length of 15 tokens.
@@ 5.3 Efficiency Evaluation
We benchmark the FlexQ W6Ax kernel across various GEMM workloads representative of different-scale LLaMA models and the OPT model. Experiments employ batch sizes of 1, 4, and 8 to assess the kernel’s adaptability to diverse inference scenarios, with particular focus on the critical GEMV operations. For comparison, we include the cuBLAS W8A8 quantized kernel and the ABQ-LLM W6A6/W6A8 kernels. Notably, due to the design of the quantization scheme, both ABQ-LLM and FlexQ employ the W6A8 kernel specifically for the ”down_proj” GEMM, which is highly sensitive to quantization accuracy.
@@
To assess FlexQ’s performance in end-to-end inference scenarios, we integrate the FlexQ W6Ax kernel into FastTransformer and compare it against multiple baselines, including the FP16 implementation of FastTransformer, the W8A16 implementation based on Cutlass, and the W8A8 version of SmoothQuant. As shown in Figure 11, results on the LLaMA-13B model demonstrate that FlexQ achieves up to 2.38× inference acceleration and 2.28× memory compression relative to FP16. Across the LLaMA and OPT model families, FlexQ consistently surpasses mainstream inference methods, delivering 1.25–1.33× speedup and 1.19–1.24× reduction in memory footprint compared to SmoothQuant. This improvement primarily stems from FlexQ’s full utilization of BTCs, which leverages the potential of 6-bit quantization to achieve higher computational efficiency.

### v2

@@ Abstract
[…]
@@ 1 Introduction
  • We perform a comprehensive evaluation on the LLaMA family and demonstrate that FlexQ achieves near-FP16 perplexity (with an increase of at most 0.1) on WikiText2, indicating negligible accuracy degradation. Our kernel significantly accelerates inference, delivering a 1.39× speedup over existing approaches like ABQ-LLM. Furthermore, end-to-end evaluations reveal that FlexQ surpasses state-of-the-art methods like SmoothQuant, delivering 1.33× faster inference and 1.21× greater memory savings.
@@ 2.1 Quantization of Large Language Models
Although LLMs have demonstrated remarkable performance, their massive size poses significant deployment challenges. Model quantization is a widely adopted technique to address these issues by producing more compact model representations. The primary focus of quantization is on the weights of linear layers (i.e., matrix multiplication), which account for approximately 99% of the total model weights (Lin et al. 2024c). Activations can also be quantized during inference to further improve efficiency (Nagel et al. 2021). Quantization schemes can be categorized into symmetric and asymmetric approaches based on whether a zero point is employed to accommodate data distribution asymmetry (Dettmers et al. 2022; Zhao et al. 2024; Liu and Yu 2025). Asymmetric quantization incorporates zero-point offsets to better capture the dynamic range of the data, thereby offering superior numerical precision. In contrast, symmetric quantization eliminates the zero-point, simplifying computations and making it more hardware-efficient. In this paper, we adopt symmetric quantization for both weights and activations. We denote the precision configuration using the notation “WxAy”, where x/y indicates the bit-width for weights and activations, respectively.
@@ 2.2 NVIDIA GPU Architecture and Tensor Core
GPU Execution Hierarchy. NVIDIA GPUs are optimized for large-scale parallel computing, comprising multiple Streaming Multiprocessors (SMs), each integrating CUDA cores, Tensor Cores, and other specialized functional units. The execution model employs a hierarchical thread organization: individual threads serve as the fundamental execution units, with 32 threads forming a warp that executes instructions synchronously in lockstep (NVIDIA 2024). Multiple warps constitute a thread block, which shares on-chip memory resources. When launching a kernel, the configuration of the number of thread blocks and threads per block must be specified. Each thread block is assigned to a specific SM for execution, where threads within the same warp operate synchronously under the Single Instruction, Multiple Threads (SIMT) execution model.
@@ 2.3 Related Work
Weight-only Quantization. Weight-only quantization approximates weight matrices with low-bit representations to reduce computational complexity and memory footprint. GPTQ (Frantar et al. 2022) utilizes 4-bit integers combined with Hessian-based error compensation to minimize quantization errors in LLMs. AWQ (Lin et al. 2024b) and OWQ (Lee et al. 2024) emphasize the importance of quantizing weights associated with higher-magnitude activations, significantly improving quantized model performance. SqueezeLLM (Kim et al. 2024) alleviates the performance degradation caused by quantization by preserving outliers and sensitive weight values within sparse matrices. QuIP (Chee et al. 2023) and QuIP# (Tseng et al. 2024) achieve 2-bit quantization by employing learnable codebooks and additional fine-tuning, often coupled with vector quantization of weights. QuaRot (Ashkboos et al. 2024) implements INT4 quantization using Hadamard transformations to effectively handle outliers in linear and attention layers. Recently, Quant-LLM (Xia et al. 2024) proposes the first GPU kernel designed specifically for FP6 weight quantization. In contrast, our work applies INT6 quantization to weights.
@@ 4.1 Quantization Scheme
Figure 4: Perplexity results (lower is better) for the LLaMA-7B and LLaMA-2-7B models, evaluating various linear layers on the WikiText2 dataset.
@@
Figure 5: Overview of the FlexQ workflow for LLaMA models, where ReQuant and DeQuant denote online quantization and dequantization operations, respectively, and BitPacking represents the online Bit-level Data Packing operation for activations.
@@ 4.3 FlexQ-W6Ax: Bit-Level Kernel Design
Specifically, for our GEMM kernel, given a weight matrix W and an activation matrix X in bit-level packed form for computation within a single thread block, we aim to optimize the memory access pattern to enable coalescing. Each thread performs load operations with a granularity of 16 bytes (128 bits). Consequently, in a single transaction, all threads (totaling threads_num) collectively transfer 16*threads_num bytes, equivalent to threads_num INT4 elements. By applying bit-level data packing to both W and X, the data can be aligned according to structures such as p*chunk_n*chunk_k or q*chunk_m*chunk_k. This alignment ensures that memory accesses are contiguous across threads, thus facilitating efficient coalesced memory transactions.
@@
Taking the mma.m8n8k128 shape as an illustrative example, the compiler partitions each warp into 8 thread groups, each comprising four consecutive threads. During data loading, the four threads within each group cooperatively load segments of the weight and activation matrices along the MMA_K dimension. Concurrently, the thread groups collaboratively handle data transfer along the [p, chunk_n] and [q, chunk_m] dimensions. Within a single iteration, each warp executes WARP_M_TILES*WARP_N_TILES BMMA operations, processing a data volume proportional to q*chunk_m*p*chunk_n*MMA_K. The intermediate results from these operations are accumulated to produce the non-weighted GEMM output.
FlexQ employs fine-grained group-wise quantization for both weights and activations. Each BMMA operation generates an intermediate result representing the multiplication of one weight group and one activation group, each associated with independently maintained quantization parameters (w_scale and x_scale). To eliminate the overhead associated with separate dequantization passes and additional memory transfers, FlexQ integrates dequantization directly into the accumulation stage of the intermediate results. Specifically, the quantization parameters w_scale and x_scale are stored as FP16 types, each thread loads the corresponding scales, which are then multiplied using the efficient half-precision parallel instruction __hmul2. The resulting dequantized intermediate values are summed across threads to produce the final FP16 result. Furthermore, by extending computations across the bit dimension among thread groups, FlexQ reduces redundant calculations, particularly in scenarios where the batch size is less than 8, thereby further optimizing GPU hardware utilization.
@@ 5.1 Experimental Setup
Dataset #Bit Method 1-7B 1-13B 1-30B 1-65B 2-7B 2-13B 2-70B 3-8B
WikiText2 FP16 - 5.68 5.09 4.10 3.53 5.47 4.88 3.31 6.14
W6A6 SmoothQuant 6.03 5.42 4.55 3.88 6.20 5.18 3.69 7.07
OmniQuant 5.96 5.28 4.38 3.75 5.87 5.14 3.71 7.24
I-LLM 5.84 5.23 4.32 - 5.68 5.10 - 6.61
DuQuant 5.74 5.13 4.15 3.60 5.53 4.92 3.35 6.27
FlexQ 5.70 5.12 4.13 3.55 5.52 4.91 3.34 6.24
C4 FP16 - 7.08 6.61 5.98 5.62 6.97 6.46 5.52 8.88
W6A6 SmoothQuant 7.47 6.97 6.34 5.99 7.76 6.76 5.88 9.57
OmniQuant 7.43 6.84 6.22 5.82 7.48 6.74 5.91 9.82
I-LLM 7.32 6.79 6.25 - 7.27 6.74 - 9.77
DuQuant 7.13 6.64 6.01 5.64 7.03 6.50 5.54 9.10
FlexQ 7.11 6.63 6.00 5.63 7.02 6.49 5.53 9.05
Table 2: Perplexity results of 6-bit weight and activation quantization applied to LLaMA family models evaluated on the WikiText2 and C4 datasets.
@@
For the accuracy experiments, we compare our method against SmoothQuant (Xiao et al. 2023), OmniQuant (Shao et al. 2024), I-LLM (Hu et al. 2024), and DuQuant (Lin et al. 2024a). Specifically, for DuQuant, we adopt the DuQuant+LWC quantization model for evaluation. To assess the efficiency of our FlexQ kernels, we compare the W6Ax variant with cuBLAS (W8A8) and ABQ-LLM (Zeng et al. 2025)’s corresponding precision kernels across various GEMM workloads on LLaMA (7B, 30B) (Touvron et al. 2023a) and LLaMA-2 (13B, 70B) models (Touvron et al. 2023b), as well as the OPT-30B model (Zhang et al. 2022). For end-to-end evaluation, we integrate our approach into FasterTransformer, comparing it against the FP16 baseline, FasterTransformer’s weight-only quantized CUTLASS (W8A16), and the SmoothQuant (W8A8) implementation. It is important to note that ABQ-LLM’s quantization strategy primarily targets extremely low-bit quantization, such as INT2. Since our focus is on INT6 quantization, we exclude ABQ-LLM from the end-to-end evaluation as a baseline.
@@
We evaluate FlexQ on multiple real-world LLMs, including LLaMA (7B–65B), LLaMA-2 (7B–70B), LLaMA-3 (8B), and OPT (6.7B–30B). Consistent with prior work (Xiao et al. 2023; Shao et al. 2024; Hu et al. 2024; Lin et al. 2024a), we report perplexity scores on language modeling benchmarks, specifically the WikiText2 (Merity et al. 2017) and C4 (Raffel et al. 2020) datasets. To evaluate zero-shot learning performance, we adopt a benchmarking methodology similar to previous studies (Frantar et al. 2022; Xiao et al. 2023; Lin et al. 2024c; Liu et al. 2025), conducting evaluations on several standard benchmarks utilizing the lm_eval harness (Gao et al. 2021), including PIQA (Bisk et al. 2020), ARC (Clark et al. 2018), BoolQ (Clark et al. 2019), HellaSwag (Zellers et al. 2019), and Winogrande (Sakaguchi et al. 2021).
@@
LLaMA3-8B
W6A6
FP16 80.85 77.78 53.41 81.28 79.16 72.84 74.22
SmoothQuant 78.94 75.88 49.49 77.58 77.39 70.80 71.68
OmniQuant 78.90 73.95 47.35 74.95 76.77 70.56 70.41
DuQuant 79.71 77.57 53.07 80.00 78.70 73.09 73.69
FlexQ 80.52 78.16 53.16 81.35 78.79 73.64 74.27
Table 4: Zero-shot common-sense QA performance of LLaMA-2 and LLaMA-3 models using 6-bit weight and activation quantization.
@@ 5.2 Accuracy Evaluation
Table 2 reports the perplexity results of FlexQ and leading baselines on the LLaMA family models. FlexQ delivers competitive results on both WikiText2 and C4, consistently outperforming SmoothQuant, OmniQuant, and I-LLM across all model sizes. Relative to the current state-of-the-art DuQuant, FlexQ is either on par or superior, trimming WikiText2 perplexity by 0.01–0.1. Remarkably, FlexQ achieves nearly identical performance to FP16 baselines while being calibration-free, offering an almost lossless and highly efficient deployment solution for LLMs.
@@
To further validate our method, we compare zero-shot accuracy of FlexQ against selected baselines across six common-sense tasks, as shown in Tables 3 and 4. FlexQ outperforms the baselines on average across all LLaMA models of different scales. Remarkably, for certain tasks, FlexQ even surpasses the accuracy of the FP16 baseline, highlighting its ability to preserve the generalization capabilities of LLM post-quantization.
@@
Table 5: Perplexity results for OPT models with 6-bit weight and activation quantization on the WikiText2 (denoted as Wiki) and C4 datasets.
@@
Figure 11: Inference latency (top) and memory utilization (bottom) measurements for the FasterTransformer implementation, evaluated on NVIDIA A6000-48GB GPU across different sequence lengths, with a comparative analysis at a fixed input length of 15 tokens.
@@ 5.3 Efficiency Evaluation
We benchmark the FlexQ W6Ax kernel across various GEMM workloads representative of different-scale LLaMA and OPT models. Experiments employ batch sizes of 1, 4, and 8 to assess the kernel’s adaptability to diverse inference scenarios, with particular focus on the critical GEMV operations. For comparison, we include the cuBLAS W8A8 quantized kernel and the ABQ-LLM W6A6/W6A8 kernels. Notably, due to the design of the quantization scheme, both ABQ-LLM and FlexQ employ the W6A8 kernel specifically for the “down_proj” GEMM, which is highly sensitive to quantization accuracy.
@@
To assess FlexQ’s performance in end-to-end inference scenarios, we integrate the FlexQ W6Ax kernel into FasterTransformer and compare it against multiple baselines, including the FP16 implementation of FasterTransformer, the W8A16 implementation based on Cutlass, and the W8A8 version of SmoothQuant. As shown in Figure 11, results on the LLaMA-13B model demonstrate that FlexQ achieves up to 2.38× inference acceleration and 2.28× memory compression relative to FP16. Across the LLaMA and OPT model families, FlexQ consistently surpasses mainstream inference methods, delivering 1.25–1.33× speedup and 1.19–1.24× reduction in memory footprint compared to SmoothQuant. This improvement primarily stems from FlexQ’s full utilization of BTCs, which leverages the potential of 6-bit quantization to achieve higher computational efficiency.